# A priori and a posteriori error estimates for discontinuous Galerkin time-discrete methods via maximal regularity

### v1

@@ 1.2. The numerical methods
${\mathcal{V}}_{k}^{\text{d}}(s):=\{v:(0,T]\to{\mathscr{D}}(A),\ v|_{J_{n}}\in{\mathbb{P}}(s),\ n=0,\dotsc,N-1\}.$
@@ 2.1. A discrete $\ell^{p}(X)$-norm and its equivalence to the continuous $L^{p}(X)$-norm
Indeed, with $\ell_{mi}$ the Lagrange polynomials $\ell_{i}\in{\mathbb{P}}_{q}$ for the Radau points $c_{1},\dotsc,c_{q},$ shifted to a subinterval $J_{m}$, we have
$\begin{split}\int_{J_{m}}\|v(t)\|_{X}^{p}\,{\mathrm{d}}t&=\int_{J_{m}}\Big{\|}\sum_{i=1}^{q}\ell_{mi}(t)v(t_{mi})\Big{\|}_{X}^{p}\,{\mathrm{d}}t\leqslant k\Big{(}\sum_{i=1}^{q}\|\ell_{mi}\|_{L^{\infty}(J_{m})}\|v(t_{mi})\|_{X}\Big{)}^{p}\\ &\leqslant\Big{(}\sum_{i=1}^{q}\|\ell_{mi}\|^{p^{\prime}}_{L^{\infty}(J_{n})}\Big{)}^{p/p^{\prime}}\Big{(}\sum_{i=1}^{q}k\|v(t_{mi})\|_{X}^{p}\Big{)}=c_{q,p}k\sum_{i=1}^{q}\|v(t_{mi})\|_{X}^{p}\end{split}$
@@
We have,
$\|v\|_{L^{p}((0,1);X)}^{p}\geqslant\int_{c_{i}-\delta}^{c_{i}}\|v(t)\|_{X}^{p}\,{\mathrm{d}}t=\int_{c_{i}-\delta}^{c_{i}}\|\ell_{i}(t)v(c_{i})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{q}\ell_{j}(t)v(c_{j})\|_{X}^{p}\,{\mathrm{d}}t,$
@@ 2.2. Maximal parabolic regularity of the dG method
see [2, Lemma 2.2]. Therefore, according to the maximal regularity result of the Radau IIA methods in [7, Lemma 3.6], we have
@@
Let us note that the estimate for the first term on the left-hand side of (2.8) is not needed in the error analysis for linear parabolic equations; however, it plays a key role in the error analysis for nonlinear parabolic equations; see [1]. The analogous estimate for Radau IIA methods from [7, Lemma 3.6] was used there in the error analysis for Radau IIA methods for nonlinear parabolic equations.
@@
Logarithmically quasi-maximal parabolic regularity results for dG methods were earlier established in [9] in general Banach spaces for autonomous equations and in [10] in Hilbert spaces for nonautonomous equations. These works cover also the cases of variable time steps as well as the critical exponents $p=1,\infty.$
@@ 2.3. An interpolant and its approximation properties
cf., e.g., [13, (12.9)] and [12, §§3.1–3.3] for the case of Hilbert spaces.
The approximation property (valid for $1\leq p\leq\infty$, $q\geq 1$)
@@ 3. Extension to nonautonomous equations
In this section, we extend the maximal parabolic regularity stability estimates for dG methods to nonautonomous parabolic equations by a perturbation argument. For similar ideas and results, we refer to [10] and [7, §3.6], [3] for the dG method with piecewise constant elements and for Radau IIA methods, respectively. Furthermore, we establish optimal order a priori and a posteriori error estimates.
@@ A.1. Existence and uniqueness of $\bm{\tilde{v}}$
see [12, (3.2)]. Indeed, the part $P_{q-2}v$ of $\tilde{v}$ on the right-hand side of (A.2) is due to the orthogonality of $v-\tilde{v}$ to the orthogonal polynomials $L_{ni},i=0,\dotsc,q-2.$ The coefficient of $L_{n,q-1}$ is determined by the interpolation condition $\tilde{v}(t_{n+1})=v(t_{n+1})$ and the property $L_{ni}(t_{n+1})=1$ of the Legendre polynomials, which yields $(P_{q-2}v)(t_{n+1})=v_{0}+\dotsb+v_{q-2}.$ In particular, $\tilde{v}=v$ for $v\in{\mathcal{X}}_{k}^{\text{d}}(q-1).$
@@
We first consider interpolation of functions on the unit interval $(0,1)$. Here we assume $1\leq p\leq\infty$ and $q\geq 1$ and hence Sobolev’s embedding ${W^{q,p}((0,1);X)}\subset C([0,1];X)$ holds. From (A.1) and (A.2) it is clear that the interpolation operator $C([0,1];X)\to{W^{q,p}((0,1);X)},v\mapsto\tilde{v},$ is bounded, so that by Sobolev’s inequality,
@@
The approximation property (A.8) for general $p$ and for Banach spaces is an analogue of [13, (12.10)] for $p=2$ and for Hilbert spaces.
@@
We begin by proving that the operator $v\mapsto\hat{\tilde{v}}$ reproduces polynomials of degree at most $q$. Again we assume $1\leq p\leq\infty$.

### v2

@@ 1.2. The numerical methods
${\mathcal{V}}_{k}^{\text{d}}(s):=\{v:(0,T]\to{\mathscr{D}}(A),\ v(0)\in X,\ v|_{J_{n}}\in{\mathbb{P}}(s),\ n=0,\dotsc,N-1\}.$
@@ 2.1. A discrete $\ell^{p}(X)$-norm and its equivalence to the continuous $L^{p}(X)$-norm
Indeed, with $\ell_{mi}$ the Lagrange polynomials $\ell_{i}\in{\mathbb{P}}_{q-1}$ for the Radau points $c_{1},\dotsc,c_{q},$ shifted to the subinterval $J_{m}$, we have
$\begin{split}\int_{J_{m}}\|v(t)\|_{X}^{p}\,{\mathrm{d}}t&=\int_{J_{m}}\Big{\|}\sum_{i=1}^{q}\ell_{mi}(t)v(t_{mi})\Big{\|}_{X}^{p}\,{\mathrm{d}}t\leqslant k\Big{(}\sum_{i=1}^{q}\|\ell_{mi}\|_{L^{\infty}(J_{m})}\|v(t_{mi})\|_{X}\Big{)}^{p}\\ &\leqslant\Big{(}\sum_{i=1}^{q}\|\ell_{mi}\|^{p^{\prime}}_{L^{\infty}(J_{m})}\Big{)}^{p/p^{\prime}}\Big{(}\sum_{i=1}^{q}k\|v(t_{mi})\|_{X}^{p}\Big{)}=c_{q,p}k\sum_{i=1}^{q}\|v(t_{mi})\|_{X}^{p}\end{split}$
@@
We have
$\|v\|_{L^{p}((0,1);X)}^{p}\geqslant\int_{c_{i}-\delta}^{c_{i}}\|v(t)\|_{X}^{p}\,{\mathrm{d}}t=\int_{c_{i}-\delta}^{c_{i}}\Big{\|}\ell_{i}(t)v(c_{i})+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{q}\ell_{j}(t)v(c_{j})\Big{\|}_{X}^{p}\,{\mathrm{d}}t,$
@@ 2.2. Maximal parabolic regularity of the dG method
see [2, Lemma 2.2]. Therefore, according to the maximal regularity result of the Radau IIA methods in [8, Lemma 3.6], we have
@@
Let us note that the estimate for the first term on the left-hand side of (2.8) is not needed in the error analysis for linear parabolic equations; however, it plays a key role in the error analysis for nonlinear parabolic equations; see [1]. The analogous estimate for Radau IIA methods from [8, Lemma 3.6] was used there in the error analysis for Radau IIA methods for nonlinear parabolic equations.
@@
Logarithmically quasi-maximal parabolic regularity results for dG methods were earlier established in [10] in general Banach spaces for autonomous equations and in [11] in Hilbert spaces for nonautonomous equations. These works are not based on the maximal regularity theory and therefore cover also the cases of variable time steps as well as the critical exponents $p=1,\infty.$
More recently, another approach to the maximal regularity of dG time discretization was presented in [6]. This is based on studying the dG approximation of a temporally regularized Green’s function. The result allows quasi-uniform meshes but is restricted to $q\geqslant 2$.
@@ 2.3. An interpolant and its approximation properties
cf., e.g., [14, (12.9)] and [13, §§3.1–3.3] for the case of Hilbert spaces.
The approximation property (valid for $1\leqslant p\leqslant\infty$, $q\geqslant 1$)
@@ 3. Extension to nonautonomous equations
In this section, we extend the maximal parabolic regularity stability estimates for dG methods to nonautonomous parabolic equations by a perturbation argument. For similar ideas and results, we refer to [11] and [8, §3.6], [3] for the dG method with piecewise constant elements and for Radau IIA methods, respectively. Furthermore, we establish optimal order a priori and a posteriori error estimates.
@@ A.1. Existence and uniqueness of $\bm{\tilde{v}}$
see [13, (3.2)]. Indeed, the part $P_{q-2}v$ of $\tilde{v}$ on the right-hand side of (A.2) is due to the orthogonality of $v-\tilde{v}$ to the orthogonal polynomials $L_{ni},i=0,\dotsc,q-2.$ The coefficient of $L_{n,q-1}$ is determined by the interpolation condition $\tilde{v}(t_{n+1})=v(t_{n+1})$ and the property $L_{ni}(t_{n+1})=1$ of the Legendre polynomials, which yields $(P_{q-2}v)(t_{n+1})=v_{0}+\dotsb+v_{q-2}.$ In particular, $\tilde{v}=v$ for $v\in{\mathcal{X}}_{k}^{\text{d}}(q-1).$
@@
We first consider interpolation of functions on the unit interval $(0,1)$. Here we assume $1\leqslant p\leqslant\infty$ and $q\geqslant 1$ and hence Sobolev’s embedding ${W^{q,p}((0,1);X)}\subset C([0,1];X)$ holds. From (A.1) and (A.2) it is clear that the interpolation operator $C([0,1];X)\to{W^{q,p}((0,1);X)},v\mapsto\tilde{v},$ is bounded, so that by Sobolev’s inequality,
@@
The approximation property (A.8) for general $p$ and for Banach spaces is an analogue of [14, (12.10)] for $p=2$ and for Hilbert spaces.
@@
We begin by proving that the operator $v\mapsto\hat{\tilde{v}}$ reproduces polynomials of degree at most $q$. Again we assume $1\leqslant p\leqslant\infty$.